OpenStudy (anonymous):

(2-3root-6)(-4+root-6)

OpenStudy (anonymous):

\[(2-3\sqrt{-6})(-4+\sqrt{-6})\]?

OpenStudy (anonymous):

try foiling first...

OpenStudy (saifoo.khan):

Lewl^

OpenStudy (anonymous):

if so rewrite as \[(2-3\sqrt{6}i)(-4+\sqrt{6}i)\]

OpenStudy (anonymous):

then multiply as you would two complex numbers \[(a+bi)(c+di)=(ac-bd)+(ad+bc)i\]

OpenStudy (anonymous):

or "foil" (ugh) you get \[(-8+18)+(12\sqrt{6}+2\sqrt{6})i\]

OpenStudy (anonymous):

aka \[10+14\sqrt{6}i\]

OpenStudy (anonymous):

i would be happy to write out the steps further if you like

OpenStudy (anonymous):

\[\sqrt[3]{125}-2\sqrt[4]{81\div16}\]